### v1

@@ Abstract
Street View-level Imagery (SVI) is a valuable data source for studies (e.g., environmental assessments, green space identification or land cover classification). While commercial SVI is available, such providers commonly restrict copying or reuse in ways necessary for research. Open SVI datasets are readily available from less restrictive sources, such as Mapillary, but due to the heterogeneity of the images, these require substantial preprocessing, filtering, and careful quality checks. We present an efficient method for automated downloading, processing, cropping, and filtering open SVI, to be used in a survey of human perceptions of the streets portrayed in these images. We demonstrate our open-source reusable SVI preparation and smartphone-friendly perception-survey software with Amsterdam (Netherlands) as the case study. Using a citizen science approach, we collected from 331 people 22,637 ratings about their perceptions for various criteria. We have published our software in a public repository for future re-use and reproducibility.
@@ 1 Introduction
Nonetheless, most urban design, transportation planning and environmental epidemiology researchers rely on objectively measured streetscape indicators because it is easier to collect measurable and quantifiable physical attributes, such as for indicators of green spaces or classification of land-cover (Kothencz and Blaschke, 2017; Xu et al., 2017; Labib et al., 2020; ElQadi et al., 2020). However, people’s subjective feelings about places do not necessarily correlate with such objectively measured spatial indicators (McCrea et al., 2006). To capture those sentiments as usable data, researchers typically use resource-intensive methods such as field interviews or questionnaires (Lynch, 1960; Lotfi and Koohsari, 2009).
Therefore, we argue that developing an efficient yet easy-to-implement method to capture human perceptions of urban environments would provide a much needed tool for answering questions like those posed above. Taken further, these answers can help with efforts to achieve the United Nations Sustainable Development Goals111www.un.org/sustainabledevelopment/sustainable-development-goals/ such as Goal 11, ‘Sustainable Cities and Communities’.
Over the past decade, various companies, organisations and people have been collecting Street View-level Imagery (SVI). The resulting datasets of SVI have been successfully used in numerous studies over a wide variety of domains (Biljecki and Ito, 2021).
Although SVI has gained momentum as a data source across disciplines, including perception collection, considerable challenges remain due to specific limitations of various SVI sources. In particular, we focus on two major classes of SVI currently accessible to researchers: (i) commercial SVI such as Google Street View (GSV), Baidu or Tencent and (ii) Volunteered Street View-level Imagery (VSVI) from open platforms such as Mapillary and KartaView (formerly OpenStreetCam), which collect geo-tagged volunteer-submitted photographs and make them available from a central repository. There are also other sources of open imagery from scientific platforms (e.g. the Global Geo-Referenced Field Photo Library) or social network sites (e.g. Flickr) that may incidentally contain useful street-level imagery for training deep learning models (Xu et al., 2017; ElQadi et al., 2020).
@@
Unlike commercial SVI platforms like Google, VSVI platforms such as Mapillary permit usage of their imagery under the Creative Commons Attribution-ShareAlike 4.0 International222creativecommons.org/licenses/by-sa/4.0/deed.en license, making it easy to adapt into an open-source software system for collecting human perception responses. However, the quality of VSVI can be vary considerably from very poor to excellent (Ma et al., 2019). Recent work by Zheng and Amemiya (2023) has highlighted ways to improve the usability of VSVI by undergoing a process of filtering, however, no studies have yet developed a comprehensive open-source toolkit to use such filtered VSVI in human perception research.
@@
Considering this gap, we aim to develop a human perception survey toolkit based on the findable, accessible, interoperable and reusable (FAIR) principles (Wilkinson et al., 2016; Barton et al., 2022) that takes a citizen science approach, which means it enables and encourages non-scientists to collaborate with scientists and contribute towards the advancement of scientific research (Fraisl et al., 2022; Haklay, 2015). This paper describes the detailed process of creating our open-source VSVI-integrated software aligned with the FAIR principles for Research Software (Hong et al., 2022), and provides an example of using this toolkit for perception data collection as a form of participatory sensing (Haklay, 2015). The survey is a mobile-friendly web application that allows participants to rate SVI based on what they see in the image. Participants may be anyone with access to a computer or a smartphone connected to the Internet. Our software presents a lightweight ‘game-like’ user interface (Bakhanova et al., 2020) that requires no more and no less than one swipe or button press per image. The source code for our survey frontend, server backend, and the VSVI filtering and pre-processing scripts, are made available under the GNU General Public License v3.0 and may be found at our Spatial Data Science and GEO AI Lab333github.com/Spatial-Data-Science-and-GEO-AI-Lab/percept web site. Together these constitute a pipeline, suitable for civic/community science (Haklay, 2015), that can be assembled by any person with basic knowledge of deploying open source software on a hosted server environment supporting ReactJS, such as virtual servers readily available for low cost from numerous providers.
@@ 2 Background
Traditional approaches to capturing human perception of space include conducting direct observations such as field surveys and resident interviews (Lynch, 1960). For instance, ‘walk-along interviews’ may be used to capture the experience of space on the spot while interviewing individuals (Carpiano, 2009; Rzotkiewicz et al., 2018). Field audits and environmental scans are two more tools for evaluating streets. For example, Van Herzele and De Vries (2012) conducted a field audit to assess the visible greenness level on the streets of two urban neighbourhoods in Gent (Belgium). Harden et al. (2024) compared the effectiveness of SVI-based ‘virtual audits’ against more traditional in-person environmental scans. Although these approaches provide a more localized assessment of human perception and lived experience of spaces, they are limited due to very high resource and time requirements. Their method does not scale to a large number of places. Considering such limitations, researchers have innovated new ways to survey human perception using massive imagery sets and a ‘crowdsourcing’ approach (Bubalo et al., 2019). We take a closer look below at several recent works that are representative of the main methods used in such research.
@@
In contrast, Twedt et al. (2016) used the direct absolute scoring approach, with a survey website where each image was shown one at a time and had to be rated on a scale of 0 to 100 using a slider adjusted with the mouse. Only 40 images were rated overall. There were approximately 300 participants, each was paid a small amount for their time via Amazon Mechanical Turk, and shown a preview of all the images before being asked to rate them individually. Pearson et al. (2024) operated similarly but scaled up to 10,727 paid participants, each of whom rated 33 images. They reported that the actual average completion time for the survey was 104 minutes, which was considerably higher than the 8 minute completion time they had expected from internal testing. Larkin et al. (2022) also paid participants through Mechanical Turk but used the Place Pulse methodology, albeit with a more precise sliding scale rating instead of the simple three choices used by the original, and they collected over 36,000 image comparisons.
Yao et al. (2019) created a human/machine feedback-loop for speeding up manual absolute scoring work while simultaneously improving the accuracy of modelled perception ratings, on the same perceptual attributes as Place Pulse (e.g. Beauty, Safety, etc).
@@
This approach sped up the manual rating process to approximately 1,000 images per hour (3.6s per image) using a recommendation algorithm based on a machine learning model. However, there are several problems with both the method and the design of the software. Firstly, one of the biggest barriers to reuse of this survey method is that it is a software application that must be downloaded and installed locally on each participant’s computer. This was not much of an obstacle for the authors because all of their participants were invited students or university staff, who were presumably comfortable with installing software from a fellow university member, but it is a major problem when trying to attract participation from the wider public. Secondly, the software’s user interface is not user-friendly because of the amount of mouse-movement required for each image: the participant adjusts a slider to a value between 1 and 100 and then they must activate a submit button (although at least some of this interaction is also possible via hot-key bindings). Reducing this user-interaction effort appears to be one of the major underlying motivations of this work. However, this leads to the third problem: to speed up the rating process, the recommendation algorithm chooses a rating that is ready to be submitted by default, but in doing so it rewards participants who simply agree with the recommendation. There is no way to distinguish true human ratings from computer-generated ratings for which the human participant did not have strong enough feelings to justify the effort of moving the mouse to change the slider away from the default setting. Therefore the algorithm risks tainting its own training data too much. Fourthly, the authors noted a significant problem, which we see as being related to (but not quite the same as) the third: the given recommendation may have the unintended consequence of influencing the participant even when they do choose to adjust the slider. For example, suppose a participant would have rated an image with a score of 20 in Beauty, but then they saw the machine recommended a score of 40, so they decided to compromise on a score of 30. There is still an element of human input in this case, but it is biased by the recommendation algorithm.
@@ 3.1 Mapillary street view imagery
We explicitly chose not to use proprietary data. For example, the standard terms of service444cloud.google.com/maps-platform/terms/ for Google SVI stresses that images are non-free and explicitly forbid the kind of usage that we need for research. In particular, in section 3.2.3 of the terms of service, they prohibit scraping, pre-fetching, bulk downloading, storing, and resharing Google SVI, all of which are necessary components of our survey method. Baidu and Tencent Maps also list several similar restrictions in their terms of service, and in any case, those platforms are mainly focused on providing imagery from China rather than worldwide.
While Google imagery is high-quality and widely available, we intend to keep within the open and FAIR principles; in particular, we are concerned that other researchers may not be able to obtain the authorization to use Google imagery within their projects (Helbich et al., 2024) and therefore we wish to show that there are alternatives that are free in the sense of being accessible and available (‘libre’) for research usage without overly-burdensome restrictions. The fact that Mapillary is free in the sense of monetary cost (‘gratis’) is also helpful. In contrast to Google, the Mapillary terms of service555www.mapillary.com/terms only require proper attribution of imagery that is downloaded and reshared or integrated into applications.
@@ 3.2 Imagery downloading and processing
Once the tile data files and photographic imagery are downloaded, we must process a very large number of JPEG files, some of which are panoramic photographs projected into a wide image format, and others which are simply plain photographs. The two main tasks we need to accomplish are: (1) finding ‘sensible subimages’ in the panoramic imagery to crop out and save separately, and (2) weeding out any
images that are too dark, too blurry or defective in some other way (such as only showing a wall or an undifferentiated block of greenery). These are both somewhat vague needs but thanks to recent advances in machine learning they can both be addressed using off-the-shelf software. In the case of panoramic photographs, we took subimages of a 4:3 ratio to show users images that fit within the app’s available view-port for imagery, rather than distorted and overly large raw panoramic images.
@@
Our definition of ‘sensible subimages’ was refined through trial and experiment, and we settled on the following: find the centers of roads in each of the panoramic images, and then for each road center crop a series of 4:3 images slightly to the left of center, on the center itself, and slightly to the right of center, as shown in Figure 2. This captures a wide variety of ways of looking at streets, including many views with a great deal of built environment and greenery, as well as views straight down the center of roads.
@@ 3.3 Filtering by image quality
An overview of all the steps of downloading and processing of Mapillary VSVI are shown in the diagram within Figure 3.
@@ 4.1 Survey set-up for the Amsterdam case study
We retrieved imagery from points found within the bounding box described by longitudes 4.7149 and 5.1220, and latitudes 52.2818 and 52.4284 (WGS 84). This encompasses the city of Amsterdam and some outlying areas. The total number of images found (panoramic and otherwise) in this bounding box was well over 700,000. The panoramic images available from Mapillary for Amsterdam are almost entirely high-resolution and professional-quality, taken from a 360-degree panoramic camera mounted on a vehicle or backpack. Using our method, each panoramic image could potentially be used to derive up to nine high-quality subimages. With so many possibilities, we heavily filtered the amount of possible SVI first by applying the pipeline as described in Section 3.3, and then selecting only images from locations closest to a fixed geographic grid of points covering the whole region but spaced approximately 20m apart from each other. This still left us with too many images, so we then down-sampled from these eligible images randomly until the number of images for the survey was a bit under 20,000, which we felt was a sufficient number for our initial run, and well-distributed across the region of interest.
@@ 4.2 Collected perception
From April 2023 to February 2024, we received 22,637 ratings across 19,750 images from up to 331 participants. We recruited participants using a wide variety of channels including but not limited to: social media, institutional mailing lists, and classrooms. We asked several socio-demographic questions of participants before they began the main part of the survey. Of those who responded, 40.5% (134) identified as women, 0.9% (3) as non-binary, 52.3% (173) as men, and the remainder declined to answer. Educational attainment amongst participants was fairly high, probably due to heavy promotion of the survey within the academic environment; 52.9% (175) people completed postgraduate education, 39.9% (132) were at the tertiary level (in the Netherlands this includes university as well as higher vocational education), 3.0% (10) had completed secondary education (high school), 0.3% (1) claimed only a primary education level and the remainder did not specify. The age spread was largely middle-aged or younger, with 48.6% (161) participants of ages from 18 to 27, 32.6% (108) aged 28 to 37, 11.2% (37) aged 38 to 47, 5.1% (17) aged 48 to 57, and the remainder older than 57. The average age of participants was about 30 (standard deviation 10). Approximately 70.7% (234) of the participants reported being based in the Netherlands, approximately 8.8% (29) reported living in another European country, a further 7.6% (25) in a non-European country, and the rest did not indicate a country or were indecipherable. About 39.0% (129) claimed a monthly gross income of less than €2,000, probably reflecting a high percentage of students, followed by about 24.2% (80) with a reported income between €2,000 and €4,000, and about 12.4% (41) claimed to have a higher income than €4,000 per month; the remainder did not specify.

### v3

@@ Abstract
Street View Imagery (SVI) is a valuable data source for studies (e.g., environmental assessments, green space identification or land cover classification). While commercial SVI is available, such providers commonly restrict copying or reuse in ways necessary for research. Open SVI datasets are readily available from less restrictive sources, such as Mapillary, but due to the heterogeneity of the images, these require substantial preprocessing, filtering, and careful quality checks. We present an efficient method for automated downloading, processing, cropping, and filtering open SVI, to be used in a survey of human perceptions of the streets portrayed in these images. We demonstrate our open-source reusable SVI preparation and smartphone-friendly perception-survey software with Amsterdam (Netherlands) as the case study. Using a citizen science approach, we collected from 331 people 22,637 ratings about their perceptions for various criteria. We have published our software in a public repository for future re-use and reproducibility.
@@ 1 Introduction
Nonetheless, most urban design, transportation planning and environmental epidemiology researchers rely on objectively measured streetscape indicators because it is easier to collect measurable and quantifiable physical attributes, such as for indicators of green spaces (Kothencz and Blaschke, 2017; Labib et al., 2020; Liu and Sevtsuk, 2024). However, people’s subjective feelings about places do not necessarily correlate with such objectively measured spatial indicators (McCrea et al., 2006). To capture those sentiments as usable data, researchers typically use resource-intensive methods such as field interviews or questionnaires (Lynch, 1960; Lotfi and Koohsari, 2009). Therefore, we argue that developing an efficient yet easy-to-implement method to capture human perceptions of urban environments would provide a much needed tool for answering questions like those posed above. Taken further, these answers can help with efforts to achieve the United Nations Sustainable Development Goals111www.un.org/sustainabledevelopment/sustainable-development-goals/ such as Goal 11, ‘Sustainable Cities and Communities’.
Over the past decade, various companies, organisations and people have been collecting Street View Imagery (SVI). Since the launch of Google Street View in 2007 (Vincent, 2007), the resulting datasets of SVI have been successfully used in numerous studies over a wide variety of domains (Biljecki and Ito, 2021). Some of the earliest research validating the usefulness of virtual streetscape audits in place of physical audits (Badland et al., 2010). Looking back at a collection of more recent work over the past decade, Dai et al. (2024, pp. 16–17) found that SVI ‘performs exceptionally well in capturing environmental variables’ and ‘holds immense potential to facilitate environmental health-related studies in the big data era’. However, we find challenges remain in the use of SVI for research purposes. In particular, we focus on two major classes of SVI currently accessible to researchers: (i) commercial SVI such as offered by Google, Bing, Baidu or Tencent and (ii) Volunteered Street View Imagery (VSVI) from open platforms such as Mapillary and KartaView (formerly OpenStreetCam), which collect geo-tagged volunteer-submitted photographs and make them available from a central repository.
@@
Unlike commercial SVI platforms like Google, which publish their content under licenses that prohibit certain research methods, VSVI platforms such as Mapillary permit usage of their imagery under the Creative Commons Attribution-ShareAlike 4.0 International222creativecommons.org/licenses/by-sa/4.0/deed.en license, making it easy to adapt into an open-source software system for collecting human perception responses. However, the quality of VSVI can be vary considerably from very poor to excellent (Ma et al., 2019). Recent work by Zheng and Amemiya (2023) has highlighted ways to improve the usability of VSVI by undergoing a process of filtering, however, no studies have yet developed a comprehensive open-source toolkit to use such filtered VSVI in human perception research.
@@
Considering this gap, we aim to develop a human perception survey toolkit based on the findable, accessible, interoperable and reusable (FAIR) principles (Wilkinson et al., 2016; Barton et al., 2022) that takes a citizen science approach, which means it enables and encourages non-scientists to collaborate with scientists and contribute towards the advancement of scientific research (Fraisl et al., 2022; Haklay, 2015). This paper describes the detailed process of creating our open-source VSVI-integrated software aligned with the FAIR principles for Research Software (Hong et al., 2022), and provides an example of using this toolkit for perception data collection as a form of participatory sensing (Haklay, 2015). The survey is a mobile-friendly web application that allows participants to rate SVI based on what they see in the image. Participants may be anyone with access to a computer or a smartphone connected to the Internet. Our software presents a lightweight ‘game-like’ user interface (Bakhanova et al., 2020) that requires no more and no less than one swipe or button press per image. The source code for our survey frontend, server backend, and the VSVI filtering and pre-processing scripts, are made available under the GNU General Public License v3.0 and may be found at our Spatial Data Science and GEO AI Lab web site (see the Software & data availability section). Together these constitute a pipeline, suitable for civic/community science (Haklay, 2015), that can be assembled by any person with basic knowledge of deploying open source software on a hosted server environment supporting ReactJS, such as virtual servers readily available for low cost from numerous providers.
@@ 2 Background
Traditional approaches to capturing human perception of space include conducting direct observations such as field surveys and resident interviews (Lynch, 1960). For instance, ‘walk-along interviews’ may be used to capture the experience of space on the spot while interviewing individuals (Carpiano, 2009; Rzotkiewicz et al., 2018). Field audits and environmental scans are two more tools for evaluating streets. For example, Van Herzele and De Vries (2012) conducted a field audit to assess the visible greenness level on the streets of two urban neighbourhoods in Gent (Belgium). Harden et al. (2024) compared the effectiveness of SVI-based ‘virtual audits’ against more traditional in-person environmental scans for assessing ‘runnability’, work that echoes one of the earliest pieces of SVI-based research: comparing virtual against physical streetscape audits for walking and cycling suitability (Badland et al., 2010). Although these approaches provide a more localized assessment of human perception and lived experience of spaces, they are limited due to very high resource and time requirements placed on individual researchers or auditors, even when the audits are conducted virtually or through photographs (Dai et al., 2024). This human-intensive method does not scale to a large number of places. Considering such limitations, researchers have innovated new ways to survey human perception using massive imagery sets (Dai et al., 2024) and a ‘crowdsourcing’ approach (Bubalo et al., 2019). We take a closer look below at several recent works that are representative of the main methods used in such research.
@@
In contrast, Twedt et al. (2016) used the direct absolute scoring approach, with a survey website where each image was shown one at a time and had to be rated on a scale of 0 to 100 using a slider adjusted with the mouse. Only 40 images were rated overall. There were approximately 300 participants, each was paid a small amount for their time via Amazon Mechanical Turk, and shown a preview of all the images before being asked to rate them individually. Pearson et al. (2024) operated similarly but scaled up to 10,727 paid participants, each of whom rated 33 images. They reported that the actual average completion time for the survey was 104 minutes, which was considerably higher than the 8 minute completion time they had expected from internal testing. Kruse et al. (2021) paid 210 Mechanical Turk workers to consider ‘playability’ on a five-point absolute scale and rated 3,011 images from 3 U.S. cities.
@@
Yao et al. (2019) created a human/machine feedback-loop for speeding up manual absolute scoring work while simultaneously improving the accuracy of modelled perception ratings, on the same perceptual attributes as Place Pulse (e.g. Beauty, Safety, etc). This approach sped up the manual rating process to approximately 1,000 images per hour (3.6s per image) using a recommendation algorithm based on a machine learning model. However, there are several problems with both the method and the design of the software. Firstly, one of the biggest barriers to reuse of this survey method is that it is a software application that must be downloaded and installed locally on each participant’s computer. This was not much of an obstacle for the authors because all of their participants were invited students or university staff, who were presumably comfortable with installing software from a fellow university member, but it is a major problem when trying to attract participation from the wider public. Secondly, the software’s user interface is not user-friendly because of the amount of mouse-movement required for each image: the participant adjusts a slider to a value between 1 and 100 and then they must activate a submit button (although at least some of this interaction is also possible via hot-key bindings). Reducing this user-interaction effort appears to be one of the major underlying motivations of this work. However, this leads to the third problem: to speed up the rating process, the recommendation algorithm chooses a rating that is ready to be submitted by default, but in doing so it rewards participants who simply agree with the recommendation. There is no way to distinguish true human ratings from computer-generated ratings for which the human participant did not have strong enough feelings to justify the effort of moving the mouse to change the slider away from the default setting. Therefore the algorithm risks tainting its own training data too much. Fourthly, the authors noted a significant problem, which we see as being related to (but not quite the same as) the third: the given recommendation may have the unintended consequence of influencing the participant even when they do choose to adjust the slider. For example, suppose a participant would have rated an image with a score of 20 in Beauty, but then they saw the machine recommended a score of 40, so they decided to compromise on a score of 30. There is still an element of human input in this case, but it is biased by the recommendation algorithm.
@@ 3.1 Mapillary street view imagery
The motivation to use Mapillary VSVI comes from the FAIR data movement (Wilkinson et al., 2016). Guidance for the final FAIR principle, reusability, emphasizes that data must be not only technically interoperable but also legally interoperable333www.go-fair.org/fair-principles/, section R1.1. However, the standard terms of service444cloud.google.com/maps-platform/terms/ for Google SVI stresses that images are non-free and explicitly forbid the kind of usage that we need for research. In particular, in section 3.2.3 of the terms of service, they prohibit scraping, pre-fetching, bulk downloading, storing, and resharing Google SVI, all of which are necessary components of our survey method. Bing, Baidu and Tencent Maps also list several similar restrictions in their terms of service, and in the case of the latter two, those platforms are mainly focused on providing imagery from China rather than worldwide.
While Google imagery is high-quality and widely available, we intend to keep within the open and FAIR principles. In particular, we are concerned that some people wishing to use our toolkit may not be able to obtain the authorization to use Google imagery, or to produce freely-reusable datasets from it, such as those who work in countries where there is legal uncertainty about these ways of using proprietary data (Helbich et al., 2024). We simply wish to show that there are alternatives that are free in the sense of being accessible and available (‘libre’) for research usage without significant restrictions on reusability. The fact that Mapillary is free in the sense of monetary cost (‘gratis’) is also helpful. In contrast to Google, the Mapillary terms of service555www.mapillary.com/terms allows usage of VSVI under a Creative Commons license that only require proper attribution of imagery that is downloaded and reshared or integrated into applications.
@@ 3.2 Imagery downloading and processing
Once the tile data files and photographic imagery are downloaded, we must process a very large number of JPEG files, some of which are panoramic photographs projected into a wide image format, and others which are simply plain photographs. The two main tasks we need to accomplish are: (1) finding ‘sensible subimages’ in the panoramic imagery to crop out and save separately, and (2) weeding out any images that are too dark, too blurry or defective in some other way (such as only showing a wall or an undifferentiated block of greenery). These are both somewhat vague needs but thanks to recent advances in machine learning they can both be addressed using off-the-shelf software. In the case of panoramic photographs, we took subimages of a 4:3 ratio to show users images that fit within the app’s available view-port for imagery, rather than distorted and overly large raw panoramic images.
@@
Our definition of ‘sensible subimages’ was refined through trial and experiment, and we settled on the following: find the centers of roads in each of the panoramic images, and then for each road center crop a series of 4:3 images slightly to the left of center, on the center itself, and slightly to the right of center, as shown in Figure 2. This captures a wide variety of ways of looking at streets, including many views with a great deal of built environment and greenery, as well as views straight down the center of roads. Kim et al. (2021) found that these choices of viewing direction, when sampling from panoramic SVI, could lead to large differences in terms of streetscape measures. Therefore, it is important for toolkit users to carefully consider the location of cropped subimages within panoramic SVI and how that might affect the results of their survey.
@@ 3.3 Filtering by image quality
An overview of all the steps of downloading and processing of Mapillary VSVI are shown in the diagram within Figure 3. After these steps, the imagery is ready to be used in the perception-gathering survey, or it may be taken and used for other purposes in compliance with Mapillary terms of service.
@@ 4.1 Survey set-up for the Amsterdam case study
The primary purpose of this paper is to introduce our SVI survey toolkit, however we conducted a case study as a demonstration that also collected some useful data. We retrieved imagery from points found within the bounding box described by longitudes 4.7149 and 5.1220, and latitudes 52.2818 and 52.4284 (WGS 84). This encompasses the city of Amsterdam and some outlying areas. The total number of images found (panoramic and otherwise) in this bounding box was well over 700,000. The panoramic images available from Mapillary for Amsterdam are almost entirely high-resolution and professional-quality, taken from a 360-degree panoramic camera mounted on a vehicle or backpack. Using our method, each panoramic image could potentially be used to derive up to nine high-quality subimages. With so many possibilities, we heavily filtered the amount of possible SVI first by applying the pipeline as described in Section 3.3, and then selecting only images from locations closest to a fixed geographic grid of points covering the whole region but spaced approximately 20m apart from each other. This still left us with too many images, so we then down-sampled from these eligible images randomly until the number of images for the survey was a bit under 20,000, which we felt was a sufficient number for our initial run, and well-distributed across the region of interest. This is effectively increasing the sparsity of sampled points, which Kim et al. (2021) found to lead to higher variability in streetscape measures; however, it was not practical to conduct our case study with significantly more images because we were not realistically expecting many tens of thousands or more responses.
@@ 4.2 Collected perception
description percentage of participants (actual number)
total participants 100% (331)
gender
     woman 40.5% (134)
     non-binary 0.9% (3)
     man 52.3% (173)
     other or not specified 6.3% (21)
education level
     postgraduate 52.9% (175)
     tertiary 39.9% (132)
     secondary 3.0% (10)
     primary 0.3% (1)
     not specified 3.9% (13)
age (average: 30, standard deviation: 10)
     18–27 48.6% (161)
     28–37 32.6% (108)
     38–47 11.2% (37)
     48–57 5.1% (17)
     58–67 1.8% (6)
     68–77 0.3% (1)
country of residence
     The Netherlands 70.7% (234)
     Another EU country 8.8% (29)
     A non-EU country 7.6% (25)
     not specified 13.0% (43)
estimated monthly gross income
     €0–1,999 39.0% (129)
     €2,000–3,999 24.2% (80)
     €4,000+ 12.4% (41)
     not specified 24.5% (81)
Table 1: Socio-demographic data summary
From April 2023 to February 2024, we received 22,637 ratings across 19,750 images from up to 331 participants. We recruited participants using a wide variety of channels including but not limited to: social media, institutional mailing lists, and classrooms. We asked several socio-demographic questions of participants before they began the main part of the survey. The responses to the socio-demographic questions are summarized in Table 1.